OpenStudy (anonymous):

Standard form for y=-4x-6

OpenStudy (anonymous):

\[y-y_1 = m(x-x_1)\] is standard form

OpenStudy (anonymous):

So it's y-y=4(x+6)?

OpenStudy (anonymous):

no. you know that the point (0,-6) exists on the line because that is the y intercept. and you know that the slope is -4. so just plug your values into the standard form equations y - (-6) = -4(x - (0)) y+6 = -4x